# Model-Independent Machine Learning Approach for Nanometric Axial Localization and Tracking

### v1

@@ Appendix A Multifocal plane optical microscopy
Conventional widefield optical microscopes have limited depth discrimination, making it difficult to accurately determine the axial position of objects, especially near the focal plane. Even with high numerical aperture objectives, a point source displaced several hundred nanometers from focus often produces nearly identical images, complicating axial localization. As shown in ref. [18], localization accuracy depends strongly on defocus: while in-focus imaging provides good precision in the transverse plane, it performs poorly along the axial axis; high defocus improves axial precision but degrades transverse accuracy. This trade-off hinders uniform 3D localization. To address this issue, the multifocal plane imaging technique was proposed in ref. [12]. This technique simultaneously captures two separate focal planes within the specimen. One focal plane matches the typical focal plane imaged by a standard wide-field microscope, whereas the other plane is offset from the standard focal plane. When the object of interest is imaged in the described setup, the image of the offset focal plane offers extra details about the object’s position. Utilization of this supplementary information enables a more precise determination of the object’s 3D location.
@@ Appendix C Optical microscope setup
The microscope [21] is equipped with a high magnification objective lens with high numerical aperture (Nikon CFI Plan Apo Lambda 100$\times$/1.45 Oil). Placing an extra magnification lens in front of the camera results in an overall magnification of 260$\times$. The critical-type custom illumination system incorporates the blue LED light source (Luminus CBT-120) with a wavelength of (460±25)460$\pm$ 25)460 ± 25 ) nm. The microscope is set up to function in reflection mode, which results in an enhanced signal-to-noise ratio. The system is equipped a fast 4-megapixel monochrome camera (Allied Vision Technologies Bonito CL-400B) operating at 100 fps. The sample is positioned within the horizontal plane using a motorized stage (Micos MS-8), and the vertical adjustment is accomplished by shifting the objective lens and the entire optical arrangement with a linear stage (Micos UPM-160). The microscope is also fitted with pneumatic vibration dampers (Fabreeka PLM 1). A Dell Precision T7500 workstation manages the microscope elements, featuring a Matrox Radient eCL framegrabber, a National Instruments PCI-7344 motion control board, and a GeForce GTX 780 GPU for enhanced image processing speed. The LASSO software framework [22, 23, 24] offers modules designed for real-time control of microscopes, featuring completely automated image capture and data analysis functionalities.

### v2

@@ Appendix A Multifocal plane optical microscopy
Conventional wide-field optical microscopes have limited axial discrimination, making it difficult to accurately determine the axial coordinate of objects, especially near the focal plane. Even with high numerical aperture objectives, a point source displaced several hundred nanometers from focus often produces nearly identical images, complicating axial localization. As shown in ref. [19], localization precision depends strongly on defocus: while in-focus imaging provides good precision in the transverse plane, it performs poorly along the axial axis; high defocus improves axial precision but degrades transverse precision. This trade-off hinders uniform 3D localization. To address this issue, the multifocal plane imaging technique was proposed in ref. [12]. This technique simultaneously captures two or more separate focal planes within the specimen. The dual-focal-plane imaging technique is a specific case in which one focal plane matches the typical focal plane imaged by a standard wide-field microscope, whereas the other plane is offset from the standard focal plane. When the object of interest is imaged in the described setup, the image of the offset focal plane offers extra details about the object’s axial coordinate. Utilization of this supplementary information enables a more precise determination of the object’s 3D location.
@@ Appendix C Optical microscope setup
The microscope [13] is equipped with a high magnification objective lens with high numerical aperture (Nikon CFI Plan Apo Lambda 100$\times$/1.45 Oil). Placing an extra magnification lens in front of the camera results in an overall magnification of 260$\times$. The critical-type custom illumination system incorporates the blue LED light source (Luminus CBT-120) with a wavelength of (460±25)460$\pm$ 25)460 ± 25 ) nm. The microscope is set up to function in reflection mode, which results in an enhanced signal-to-noise ratio. The system is equipped a fast 4-megapixel monochrome camera (Allied Vision Technologies Bonito CL-400B) operating at 100 fps. The sample is positioned within the horizontal plane using a motorized stage (Micos MS-8), and the axial adjustment is accomplished by shifting the objective lens and the entire optical arrangement with a linear stage (Micos UPM-160). The microscope is also fitted with pneumatic vibration dampers (Fabreeka PLM 1). A Dell Precision T7500 workstation manages the microscope elements, featuring a Matrox Radient eCL framegrabber, a National Instruments PCI-7344 motion control board, and a GeForce GTX 780 GPU for enhanced image processing speed. The LASSO software framework [22, 23, 24] offers modules designed for real-time control of microscopes, featuring completely automated image capture and data analysis functionalities.